# Large Language Models in Mental Health Care: a Scoping Review

### v1

@@ 3.4 Validation measures and metrics
The chatbot’s ability to keep the conversation focused and transition smoothly between topics.
Language
3.5 Ethics, Privacy, Safety, and Other Concerns
Privacy is a major concern across studies, particularly amplified due to the sensitive nature of data often processed by mental health care applications. Several studies underscore this risk of sensitive data exposure and stress the need for robust data protection and ethical standards.[44, 66, 45, 62] Safety and reliability are identified as foundational requirements, with an emphasis on preventing harmful content generation [19, 49, 61, 62] and ensuring accurate and appropriate responses.[119, 45, 74, 67] The literature also draws attention to the essential balance between leveraging LLM’s benefits and ensuring safety, with a focus on risk assessment in mental health support [44, 59, 47, 51] and a continuous pursuit of reliability and consistency.[119, 48, 50, 61]
An increasing dependency on AI for emotional support or therapeutic guidance is a notable concern,[19, 73, 46, 49] raising issues about the neglect of real-life interactions and the potential for over-reliance. This is exacerbated by the implications of inaccurate or biased content, which can seriously affect perceptions and decisions in mental health contexts.[44, 65, 67, 54] Anthropomorphism is pinpointed as an ethical consideration, influencing user expectations and reliance on AI systems.[19] The intimate nature of mental health discussions amplifies the risk of misunderstandings and misuse,[66, 60, 67, 62] underscoring the importance of responsible LLM usage.
Technical and performance challenges are also frequently mentioned, ranging from model limitations and generalizability [44, 119, 47, 55] to memory and context limitations.[19, 66, 50, 62] These issues impact the reliability and efficacy of AI applications, particularly in complex, real-world settings. Variability in performance [59, 46, 54] and the need for robustness and transparency [59, 51, 61, 63] are areas that require continuous innovation and scrutiny.
The transition to real-world applications brings additional complexity, particularly in mental health where accuracy and sensitivity are crucial. Discussions on long-term effectiveness,[73, 47, 76, 58] the distinction between controlled and real-world settings,[45, 48, 51, 62] and addressing the accessibility and digital divide [46, 52, 63, 58] illustrate the challenges in bridging the gap between LLMs’ potential and its practical implementation.
Finally, a few studies noted the need for diverse and large datasets,[65, 60, 67, 62] specialized training,[119, 59, 46, 63] and consideration of data annotations.[66, 67, 62, 76] The necessity of substantial computational resources [44, 50, 51, 62] and expert involvement [66, 47, 61, 62] are also highlighted as critical components in advancing the field responsibly.
@@ 3.6 Benchmark studies
Benchmark studies are included in this application-focused review to identify areas needing enhancement for objectively assessing the efficacy of various LLMs in mental health care. So far there have been two benchmark studies.[119, 67] Jin et al. [119] provides a comprehensive evaluation across models including GPT-4, GPT-3.5, Alpaca,[79] Vicuna,[80] and LLaMA-2,[37] on six tasks: diagnosis prediction, sentiment analysis for inferring emotional states, language modeling for understanding conversations, and question answering for therapeutic guidance, along with evaluating therapy session dialogues. Data from social media and therapy sessions. In contrast, Qi et al. [67] have a much more focused scope on classifying cognitive distortions and predicting suicidal risks on Chinese social media data. They evaluated ChatGLM2-6B,[129] GLM-130B,[130] GPT-3.5, and GPT-4.
@@ 4 Discussion
This study is the first review of Large Language Models’ (LLMs) progression in mental health care since the T5 model’s inception in 2019. Through an examination of 34 pertinent studies, we synthesized the characteristics, methodologies, datasets, validation measures, application areas, and specific mental health challenges addressed by LLMs. This synthesis aims to serve as a bridge between the computational and mental health communities, detailing insights in an accessible manner.
LLMs, as a class of algorithms, have demonstrated remarkable capabilities in NLP. Their skillset aligns closely with the requirements for mental health-related tasks, indicating their potential as foundational tools in this field. However, despite preliminary evidence suggesting the feasibility of LLMs in enhancing mental health care, notable gaps remain between their current state and actual clinical applicability. Below, without being exhaustive, we outline three key areas for improvement that could help realize the full potential of LLMs in clinical practice.
@@
Current fine-tuning efforts mainly utilize dialogue and social media data, like Reddit postings. While this data can be useful for screening, it often falls short of establishing formal clinical diagnoses due to oversimplifications and the diverse nature of clinical symptoms.[131, 132] In addition, they were sourced from a distribution of users that very likely do not represent the general population at risk of the mental conditions in question, and descriptions of compositions of the included users are lacking. Moreover, labels in existing datasets are typically not curated or validated by mental health experts, with substantial variability in their definitions across studies. For example, "depression" and "anxiety" range from everyday mood fluctuations to clinical disorders. Only one study [65] made a distinction between clinical diagnoses and self-reported mental states. For LLMs to be reliably integrated into clinical workflows, development and validation with high-quality data are crucial. This data should accurately define mental health conditions and provide a solid basis for rigorous assessment.
Enhanced reasoning, empathy capabilities, and evaluation methods: Effective dialogue-based tasks, especially those pertaining to mental health care, require advanced reasoning and empathy abilities, allowing for the analysis of statements provided by the users, and the provision of appropriate feedback. However, the extent to which LLMs’ reasoning is a product of their large training datasets or a genuine emergent property remains a subject of debate (i.e., memorization vs. understanding/reasoning). In cases where "reasoning" is based on memorized training data, the models may lack robustness, particularly in handling diverse user dialogues and language use. In addition, although LLMs have shown preliminary capabilities in demonstrating empathy, these results are still inconclusive.
The absence of a unified evaluation framework amplifies this issue. While discriminative tasks with clear right or wrong answers are easy to assess, evaluating free-text generation is challenging due to the absence of a universal standard. Previous studies have used automated metrics like BLEU, ROUGE-L, and Meteor to evaluate language quality, but these fail to capture context. Even advanced metrics like the BERT score and GPT-3 score fall short in capturing the nuances of mental health care dialogues. Most studies have resorted to human evaluations, but the lack of a standardized framework leads to subjective selection of assessment aspects, resulting in varied and potentially incomparable results. This methodology discrepancy hinders the generalization of findings and underscores the need for a more relevant evaluation framework. Future research should also focus on developing a standardized validation metric and framework for assessing LLMs’ free-text generations in mental health care.
@@ 5 Conclusion
LLMs hold promise for enhancing mental health care with their ability to process and generate large volumes of textual information. Despite their potential, considerable challenges in data quality, model reasoning, empathy, and evaluation methods must be addressed. Moreover, ensuring privacy, safety, and ethical standards is crucial for their integration into clinical settings. Future advancements depend on collaborative efforts to refine technology, develop standardized evaluations, and ensure ethical applications, aiming to realize LLMs’ full potential in supporting mental health care.

### v3

@@ 3.4 Validation measures and metrics
Coherence [117, 46, 48] The chatbot’s ability to keep the conversation focused and transition smoothly between topics. Language
Consistency [56, 48] Ensuring the chatbot’s responses are aligned with its defined role without contradictions. Language
Empathy [40, 45, 49, 50, 53] The degree to which the chatbot shows warmth, sympathy, and concern towards the user’s situation. Mental Health
Engagement [69, 49, 50] The chatbot’s ability to maintain prolonged and engaging conversations with the user. Mental Health
Exploration [117, 41, 69, 47] Indicators of the user’s attempts to investigate deeper into the topics brought up in the conversation. Mental Health
Fluency [40, 117, 55, 42, 46, 48] The chatbot’s generation is grammatically correct, well-formed, and natural-sounding, resembling human-like speech or writing. Language
Guidance [43, 44] The chatbot’s effectiveness in giving clear, actionable instructions and guidance. Mental Health
Helpfulness [42, 48, 50, 53] The extent to which the chatbot contributes to reducing the user’s emotional distress and improving mood. Mental Health
Informativeness [43, 48] The extent to which the chatbot provides useful and relevant suggestions for the user’s issues. Mental Health
Interpretation [117, 68, 43, 44, 47] The user’s effort to restate or reinterpret the problems presented, as indicated by linguistic markers. Mental Health
Professionalism [56, 44, 46] Evaluating the chatbot’s responses for their logical and psychological accuracy. Mental Health
Relevance [40, 117, 42, 43, 70] The chatbot’s ability to generate content related to the user’s specific problems. Mental Health
Understanding [48, 49] The chatbot’s level of accurately grasping and reflecting the user’s experiences and emotions. Mental Health
Reliability [55, 56] The chatbot’s generation is factually accurate, logically consistent, and dependable across similar inputs or contexts. Mental Health
@@ 3.5 Benchmark studies
Benchmark studies are included in this application-focused review to identify areas needing enhancement for objectively assessing the efficacy of various LLMs in mental health care. So far there have been two benchmark studies[117, 62]. Jin et al. [117] provides a comprehensive evaluation across models including GPT-4, GPT-3.5, Alpaca[75], Vicuna[77], and LLaMA-2[33], on six tasks using data from social media and therapy sessions: diagnosis prediction, sentiment analysis for inferring emotional states, language modeling for understanding conversations, and question answering for therapeutic guidance, along with evaluating therapy session dialogues. In contrast, Qi et al. [62] have a much more focused scope on classifying cognitive distortions and predicting suicidal risks on Chinese social media data. They evaluated ChatGLM2-6B[126], GLM-130B[127], GPT-3.5, and GPT-4.
@@ 4 Discussion
Through an examination of 34 pertinent studies, we synthesized the characteristics, methodologies, datasets, validation measures, application areas, and specific mental health challenges addressed by LLMs. This synthesis aims to serve as a bridge between the computational and mental health communities, detailing insights in an accessible manner. Despite preliminary evidence suggesting the feasibility of LLMs in enhancing mental health care, our results show that notable gaps remain between their current state and actual clinical applicability. Below, we outline three key areas for improvement that could help realize the full potential of LLMs in clinical practice.
@@
Current fine-tuning efforts mainly utilize dialogue and social media data, like Reddit postings. While this data can be useful for screening, it falls short of establishing formal clinical diagnoses due to oversimplifications and the diverse nature of clinical symptoms[128, 129]. In addition, they were sourced from a distribution of users that very likely do not represent the general population at risk of the mental conditions in question, and descriptions of compositions of the included users are lacking. Moreover, labels in existing datasets are typically not curated or validated by mental health experts, with substantial variability in their definitions across studies. For example, ”depression” and ”anxiety” range from everyday mood fluctuations to clinical disorders. Only one study [60] made a distinction between clinical diagnoses and self-reported mental states. For LLMs to be reliably integrated into clinical workflows, development and validation with high-quality data are crucial. This data should accurately define mental health conditions and provide a solid basis for rigorous assessment.
Enhanced reasoning, empathy capabilities, and evaluation methods: Effective dialogue-based tasks, especially those on mental health care, require advanced reasoning and empathy abilities, allowing for the analysis of statements provided by the users, and the provision of appropriate feedback. However, the extent to which LLMs’ reasoning is a product of their large training datasets or a genuine emergent property remains a subject of debate (i.e., memorization vs. understanding/reasoning). In cases where ”reasoning” is based on memorized training data, the models may lack robustness, particularly in handling diverse user dialogues and language use. In addition, although LLMs have shown preliminary capabilities in demonstrating empathy, these results are still inconclusive.
The absence of a unified evaluation framework amplifies this issue. While discriminative tasks with clear right or wrong answers are easy to assess, evaluating free-text generation is challenging due to the absence of a universal standard. Previous studies have used automated metrics like BLEU, ROUGE-L, and Meteor to evaluate language quality, but these fail to capture context. Even advanced metrics like the BERT score and GPT-3 score fall short in capturing the nuances of mental health care dialogues. Most studies have resorted to human evaluations, but the lack of a standardized framework leads to subjective selection of assessment aspects, resulting in varied and potentially incomparable results. This methodology discrepancy hinders the generalization of findings and underscores the need for a more relevant evaluation framework.
@@
Continuing Trends in 2024: The ongoing scarcity of robust clinical datasets in mental health research likely sustains heavy reliance on social media-derived data (e.g., Reddit/X) as primary sources for model development in 2024. This trend mirrors our documented observations of their pervasive use in 2023 for training and validating mental health-related AI tools, raising critical questions about inherent biases, representativeness, and clinical generalizability.
The absence of standardized validation frameworks persists as a key methodological challenge. Our analysis identified marked heterogeneity in evaluation protocols during 2023, with no consensus emerging on performance metrics or clinical benchmarking criteria. This inconsistency will likely have extended into 2024, underscoring an enduring barrier to rigorous assessment of LLM efficacy, safety, and translational potential in mental healthcare contexts.
Furthermore, accelerating interdisciplinary integration between AI developers, clinical practitioners, and public health stakeholders, a trend catalyzed in 2023 by proliferating open-source LLM ecosystems, has probably intensified in 2024. Such collaborations hold promise for advancing context-sensitive model architectures, yet they concurrently necessitate robust governance frameworks to address ethical risks and ensure equitable implementation.
@@ 5 Conclusion
LLMs hold promise for enhancing mental health care with their ability to process and generate large volumes of textual information. Our results suggest that while it is feasible to run these models, concerns about the lack of valid training sets, lack of standardized evaluation metrics, and privacy/ethical concerns require careful consideration before any clinical implementation. Addressing these challenges will be more productive than further pilot studies which show the feasibility but also require new collaborations to refine technology, develop standardized evaluations, and ensure ethical applications, aiming to realize LLMs’ full potential in supporting mental health care.
Key References
This section highlights several key references published within the review period that exemplify or directly inform the major challenges and future directions outlined in the Discussion, particularly concerning data quality, model capabilities, evaluation, and ethical considerations.
  • Xu et al. [60]: As noted in the discussion, this work is important for attempting to differentiate between clinical diagnoses and self-reported mental states, directly addressing the crucial challenge of data definition and quality for reliable LLM application in mental health.
  • Harrigian et al. [128]: This study, cited in the discussion, critically examines the generalizability of mental health models trained on social media data, underscoring the need for improved, clinically relevant datasets emphasized in this review.
  • Lee et al. [47]: This paper exemplifies recent efforts to enhance LLM capabilities crucial for mental health care, specifically empathy. It directly addresses the need for improved model functionalities and nuanced interaction methods discussed as vital for progress, even if current results remain inconclusive.
  • Jin et al. [117]: Directly confronts the lack of standardized evaluation metrics identified as a major hindrance in the discussion. Proposing a dedicated benchmark for mental health LLMs represents a critical step towards comparable assessments of model performance and safety.
  • Yang et al. [56]: This study exemplifies the fine-tuning of open-source models (LLaMA 2) discussed as necessary to move beyond closed models like GPT-4. It also showcases work directly using social media data, highlighting both the potential and the challenges related to data quality and interpretation raised in the discussion.
  • Ma et al. [115]: This study provides user perspectives on LLM-based mental health support, directly informing the discussion around practical implementation challenges, including ethical considerations, safety, and the potential for over-reliance.
Appendix A Statements and Declarations
A.1 Funding
The authors declare that no funds, grants, or other support were received during the preparation of this manuscript.
A.2 Competing Interests
JT has research support from Otsuka and is an adviser to Precision Mental Wellness. All other authors have no conflict of interest.
A.3 Author Contributions
Study design, data collection, and article screening: Y.H.
Result analysis: Y.H., F.L., K.Y., and Z.L.
Manuscript drafting: Y.H., F.L., K.Y., Z.L., H.N., P.Z, and Y.S.
Critical feedback: Y.S., L.M., and A.B.
Manuscript review: All authors.
Y.H. takes the integral responsibility of the study.
Appendix B Supplementary
B.1 S.1. Screening prompt for ChatGPT-4
The template shown in Figure 3 was employed for screening titles and abstracts in the review’s screening phase. After testing various prompts, the one presented here demonstrated the highest accuracy, achieving a perfect score (10 out of 10) in a randomly selected sample. It was crucial to specify that articles should be explicitly focused on LLMs and mental health care. This precision was necessary to avoid the inclusion of articles that were only vaguely related to the topic, a common issue encountered when the criteria were less specific.
Figure 3: Screening prompt for ChatGPT-4.
S.2.: Data Restriction and Licensing in Related Datasets
Dataset restrictions and licenses establish the guidelines for how data may be utilized, shared, altered, and disseminated. Users must comply with the terms set forth by the restrictions and licenses under which the datasets are released. Below is an explanation of the various restrictions and licenses associated with the datasets we reviewed:
  1. 1.
    Public Available: Indicates that the dataset is made freely available to the public and can be used, distributed, or modified without any restrictions, sometimes even for commercial purposes, depending on the exact nature of the public domain status.
  2. 2.
    Research Only: Indicates that the dataset is available only for non-commercial, academic research purposes. It typically restricts commercial use or widespread redistribution without explicit permission.
  3. 3.
    MIT License: A permissive free software license originating at the Massachusetts Institute of Technology (MIT). It allows for almost unrestricted freedom with the work, including commercial use, with the only requirement being to preserve the copyright and license notices.
  4. 4.
    Apache 2.0: A permissive free software license written by the Apache Software Foundation. It allows users to use the software for any purpose, to distribute it, to modify it, and to distribute modified versions of the software under the terms of the license, without concern for royalties.
  5. 5.
    CC-BY-4.0 (Creative Commons Attribution 4.0 International): Allows others to share, copy, distribute, execute, and transmit the work, as well as to remix, adapt, and build upon the work, even commercially, as long as they provide a credit to the original creator.
  6. 6.
    CC-BY-NC-4.0 (Creative Commons Attribution NonCommercial 4.0 International): Similar to CC-BY, this license allows others to remix, adapt, and build upon the work non-commercially, and although their new works must also acknowledge the creator and be non-commercial, they don’t have to license their derivative works on the same terms.
Datasets that do not have a specific license mentioned or that the availability of the dataset is not explicitly stated are marked with N/P. It is advisable to contact the dataset providers or creators for more information on usage rights and restrictions